# Optimal Execution in Intraday Energy Markets under Hawkes Processes with Transient Impact

### v1

@@ 3.1 Estimation of the Hawkes Parameters
We estimate the Hawkes process parameters using maximum likelihood estimation (MLE), which is established in the literature (e.g., [32]). The log-likelihood for multivariate Hawkes processes, along with its gradient and Hessian, is detailed in Appendix A. Numerical optimization is performed using a quasi-Newton method.
@@
For the baseline intensity, we employ a non-parametric approach using cubic B-splines and cosine bump functions. B-splines are particularly effective for modeling time-varying intensity patterns, such as the U-shape observed in equity markets [13, 11]. To our knowledge, this non-parametric framework has not been applied to intraday energy markets.
The baseline intensity is defined as:
@@
where $i=1,2,\dots,N_{\text{basis}}$, $w=\frac{l}{N_{\text{basis}}-3}$, and $l=T_{e,m}-T_{b,m}$. Here, $N_{\text{basis}}$ represents the number of basis functions, which can be determined using a grid search.
While increasing $N_{\text{basis}}$ improves the model’s flexibility, it also increases computational complexity by enlarging the parameter space, making the non-convex optimization problem more challenging.
@@
Furthermore, Figure 5 highlights that the median calibrated spline functions capture the slight seasonality across trading products. Specifically, the intensities exhibit an increase from the early hourly products, peak around the mid-day products, and slightly decrease for the afternoon products. These patterns align with the empirical analysis in Section 1.1, particularly the observations illustrated in Figure 2 for the buy-side order book (LOB).
@@
Figure 6: Empirical and fitted intensity for Buy/Sell Market order. We estimate the empirical intensity by counting the average number of events during $[t,t+\delta t]$, where $N(t)$ the realized counting process for each side of the LOB.
@@
The median branching ratio, defined as $\frac{\alpha}{\beta}$ in the univariate case, represents the average number of arrivals triggered by the excitation from a single event. For the spline approximation, this ratio is approximately $0.30$ to $0.32$ for most calibrated products on the Buy side, satisfying the stationarity condition $\frac{\alpha}{\beta}<1$ required for univariate exponential Hawkes processes [6]. This implies that the arrival of a Buy market order triggers, on average, an additional $0.30$ to $0.32$ Buy market orders.
Kolmogorov-Smirnov (KS) tests reveal no significant differences at the $5\%$ level between the Buy and Sell sides in the distributions of excitation parameters and branching ratios. Additionally, no distinct patterns are observed in the calibrated parameters of the exponential excitation kernel across different delivery hours. However, the calibrated baseline intensity using the spline approximation exhibits noticeable variations across products (see Figure 5), effectively capturing the observed increase in market activity for mid-day traded products, as illustrated in Figure 2.
@@
Table 1: Medians ($\pm$ std) of the optimal parameters of the self-excitation exponential kernel $t\mapsto\alpha\mathop{}\!\textnormal{e}^{-\beta t}$ for Buy and Sell market order arrivals conditionally on the delivery hour, medians of branching ratios and half-lives.
@@
Figures 7 and 8 present QQ-plots of inter-event times for all events on the bid side of the order book, focusing on two selected maturities (19h and 21h). For short inter-event times ($\Delta\tau\leq 300s$), the model provides a good fit. However, the fit deteriorates for longer inter-event times, particularly for the shorter maturity (19h). This discrepancy may be attributed to lower trading activity at the start of the trading session. In some sessions, this inactivity significantly extends the tail of the inter-event time distribution, making it slightly left-skewed. This behavior is not fully captured by the calibrated model, potentially limiting its ability to accurately simulate the distribution of long inter-event times.
Figure 7: QQ-plot for Buy MOs (Simulated vs Realized) for maturity 21h
Figure 8: QQ-plot for Buy MOs (Simulated vs Realized) for maturity 19h
@@
So far we have considered only the case where the cross-covariances are equal to zero and whilst the counting process $\left(N_{t}^{+},N_{t}^{-}\right)$ describing the arrival of Buy/Sell market orders respectively may be self-exciting, they are independent and thus not mutually exciting. To introduce the excitation between the two types of orders we consider the parameterization of the conditional intensity function given by the full impact matrix kernel (17). Following the Goodness of Fit analysis we have conducted (See Appendix B.1) in the univariate case we can see that overall the fitted model with the spline approximation of the baseline intensity provides a better fit to the data. Hence as a baseline intensity, we will consider the one given by the spline approximation. We present the calibrated parameters in the bi-variate case below. Our goal is to test the performance of the two models to cost savings compared to the benchmark models; this will be considered in the following section.
In order to calculate the resilience speed of the transient part of the price we will use the approximation derived in [14]. The authors derived the resilience speed parameter for a TWAP type of schedule, assuming that the market orderflow follows a self-exciting Hawkes process with exponential kernel $t\mapsto\alpha\mathop{}\!\textnormal{e}^{-\beta t}$. Under this setting they showed that a good approximation of the resilience speed is $\rho\approx\frac{1-\alpha/\beta}{T/2}$, where $[0,T]$ the time period where the trader want to liquidate his position. The approximation shows how the resilience parameter links to the self-exciting dynamics of market order flow. As the authors mention this resilience speed compromises only an approximation since different liquidating strategies would in principle lead to different resilience speed parameters $\rho$. To this end we follow this approach and we leave out the calibration of a propagator type of model for future research.
@@ 3.2 Execution Costs
We will now present a summary of the cost estimates for the different liquidation strategies we are going to propose in the following section. In order to perform such a transaction cost analysis we should first find some appropriate cost estimates. Since we do not have any endogenous data, we proceed to estimate various market impact factors from public data disseminated by the exchange and the LOB [21]. Our calibration methodology is similar in nature to the analysis presented in [11, 23] where an approximation of the half-spread and instantaneous trading costs is studied using LOB market data.
@@ 3.2.1 Instantaneous impact estimation from stochastic order book
Thus, the LOB state at any given time can be fully described by the random vector $\{(M_{i},\Delta P_{i})\}_{i=0,\dots,K_{0}}$. Compared with simple snapshot averages, a non-parametric estimation of the joint distribution $\{(M_{i},\Delta P_{i})\}$ can capture more nuanced stochastic behavior—including asymmetries and heavier tails—leading to better estimates of transaction costs.
@@
$P(\nu):=\frac{1}{\nu}\,\mathbb{E}\bigg{[}$ $\sum_{j=1}^{K}\mathbf{1}_{\{m_{j-1}<\nu\leq m_{j}\}}\Big{(}\sum_{i=1}^{j}M_{i-1}\,\Delta P_{i-1}$ (27)
$+(\nu-m_{j})\,\Delta P_{j}\Big{)}\bigg{]},$
where $\nu$ is the size of a large market order. In general, estimating such a high-dimensional joint distribution is prone to the curse of dimensionality, requiring a balance between bias and variance. To alleviate complexity, a modified cost function can be used to reduce dimensionality while still offering a meaningful estimate of the instantaneous impact.
@@
$\hat{P}(\nu)$ $:=\frac{1}{\nu}\sum_{j=1}^{K}\mathbb{E}\Bigl{[}M_{j-1}\,\Delta P_{j-1}+(\nu-m_{j})\,\Delta P_{j}\;\Big{|}\;V_{i,j}\Bigr{]}$
×$\qquad\times\mathbb{P}\big{(}m_{j}<\nu\leq m_{j+1}\big{)},$ (28)
where $V_{i,j}=\{\omega\in\Omega\colon m_{j}<\nu\leq m_{j+1}\}$ and $\nu$ is the market order size, $\Delta P_{j}$ are the price offsets, and $M_{j}$ denotes the liquidity at offset $\Delta P_{j}$. The index $m_{j}$ is the aggregate volume up to level $j$.
@@
To approximate the joint density of the random variables $x=m_{i}$ and $y=m_{i+1}$, we use a bivariate kernel density estimator. Suppose we have $\bigl{\{}(x_{t},y_{t})\bigr{\}}_{t=1}^{n}$ drawn from the distribution of $(m_{i},m_{i+1})$. The estimator is:
@@
To estimate the instantaneous market impact, we shall simulate he order book in its stationary state, followed by the submission of orders of size $\nu$. For each order, we compute $\hat{P}(\nu)$ from (3.2.1). Next, we fit a linear relationship:
$\hat{P}(\nu)=\eta\,\nu+\varepsilon,\quad\varepsilon\sim\mathcal{N}\bigl{(}0,\sigma^{2}\bigr{)},$
where $\hat{P}(\nu)$ is the price change per MWh after executing an order of size $\nu$ MWh. We interpret $\eta$ as the instantaneous market impact factor.
@@
Figure 9 shows the average piecewise-interpolated function describing the linear execution cost coefficient. The median calibrated $\hat{\eta}(t)$ for each hourly contract clearly varies over time; specifically, $\hat{\eta}(t)$ decreases after the trading session begins (measured relative to the last eight trading hours) and slightly increases in the final hour (local trading regime). This aligns with the notion that liquidity diminishes near gate closure, causing higher transaction costs [8, 23, 26].
Figure 9: Estimated temporary impact coefficients with respect to gate closure for the sell side of the LOB. These impact coefficients are obtained by simulating a large market order that “walks” through the limit order book, assuming a linear relationship $\hat{P}(\nu)=\eta\,\nu+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$ and $\hat{P}(\nu)$ denotes the price change per MWh for an order of size $\nu$.
@@ 3.2.2 Bid-Ask Spread Transaction Costs
Following [8, 23, 26], the bid-ask spread (BAS) in intraday energy markets exhibits a clear seasonal pattern. In particular, the average BAS decreases significantly at the outset of the trading session and remains nearly constant for the final eight hours of trading. One hour before maturity, when trading is restricted to the local market, the spread widens sharply. This behavior mirrors the pattern observed for the half-spread, which closely tracks the shape of the temporary impact coefficient $\hat{\eta}(t)$. Both observations suggest that liquidity is notably more limited at market open and after market decoupling compared to just before decoupling [23, 26].
Consequently, the cost of immediacy, i.e. submitting a market order instead of a limit order—tends to be substantially higher at the start of trading and during the final trading hour. In contrast, during the intermediate trading period, these costs remain relatively lower (see Figure 9).
@@ 4.1 Scaling and discretization of the optimal strategy
Before describing the methodology followed for the back testing of the optimal strategy, we will first describe the simplest and most natural way to use the optimal strategy described in the equation (4.1). As we have already presented the optimal strategy can be written as
$X_{t}^{*}$ $=-(1-\epsilon)^{-1}(1+\rho hD_{t}^{*}$
$\quad+\frac{m_{1}}{2\rho}\left(2+\rho h\left(1+\zeta(\eta h+\mu\rho h\omega(\eta h\right)\right)\delta_{t}.$ (31)
We recall that $\delta_{t}=\lambda_{t}^{+}-\lambda_{t}^{-}$, is the vector of intensity imbalances and are modeled with one of the aforementioned models using the seasonal versions of the baseline intensity, which has been estimated non-parametrically as described in Section 3. We are then going to consider the univariate and bi-variate models separately for the backtesting and transaction cost analysis study using the calibrated parameters obtained (See Table 1 for the univariate case and Table 5 for the bi-variate case). In order to use the optimal liquidating strategy we consider a subset $\Theta$ of the time interval $\Theta\subset[T_{b,m},T_{e,m}]\subset[0,T]$ possible made of stopping times and trade for each $t\in\Theta$ the quantity
@@
$+\frac{m_{1}}{2\rho}\times\left[1+\frac{\rho h}{2+\rho h}\times\left[\zeta h\eta)+\mu\rho\omega(h\eta)\right]\cdot s\delta_{t}\right],$
so that equation (4.1) holds for $t+$ if $s=1$. In order to calculate the resilience part of the execution price we consider the simplified expression
$D_{t}^{*}=\sum_{\tau\leq t}\Delta N_{\tau}\left[G(t-\tau)-G(\infty)\right],$ (33)
The propagator kernel is chosen as in the OW framework $G(u)=(1-\mu)\exp(-\rho u)+\mu$ and $\Delta N_{\tau}$ the jumps of the order-flow. As we have already described we have a positive (negative) unit jump when we have an arrival of a buy (sell) market order. We also introduce the scaling factor $s\in[0,1]$ that multiplies both $\delta_{t}$ and $D_{t}$ and is a hyper-parameter that tunes the leveraging effect of the strategy, describing the deviation of the whole strategy from the standard liquidation scheme in [31]. We can observe that in the case $s=0$ we derive the static strategy, which is nothing else by the block trades that have been derived in the OW strategy, and in the case $s=1$ we derive the strategy given by optimal execution strategy we derived under the Marked Hawkes process which drives the order flow and presented in Theorem 2.2.

### v2

@@ 3.1 Estimation of the Hawkes Parameters
We estimate the Hawkes process parameters using maximum likelihood estimation (MLE), which is established in the literature (e.g., [Ozaki1979MaximumLE]). The log-likelihood for multivariate Hawkes processes, along with its gradient and Hessian, is detailed in Appendix A. Numerical optimization is performed using a quasi-Newton method.
@@
For the baseline intensity, we employ a non-parametric approach using cubic B-splines and cosine bump functions. B-splines are particularly effective for modeling time-varying intensity patterns, such as the U-shape observed in equity markets [Chen_Hall_2013, cartea2015algorithmic]. To our knowledge, this non-parametric framework has not been applied to intraday energy markets. The baseline intensity is defined as:
@@
where $i=1,2,\dots,N_{\text{basis}}$, $w=\frac{l}{N_{\text{basis}}-3}$, and $l=T_{e,m}-T_{b,m}$. Here, $N_{\text{basis}}$ represents the number of basis functions, which can be determined using a grid search. This construction provides the flexibility to reproduce the empirical shape of market activity given by an exponential rise toward gate closure followed by a sharp decline, while retaining smoothness and interpretability. While increasing $N_{\text{basis}}$ improves the model’s flexibility, it also increases computational complexity by enlarging the parameter space, making the non-convex optimization problem more challenging.
@@
Furthermore, Figure 4 highlights that the median calibrated spline functions capture the slight seasonality across trading products. Specifically, the intensities exhibit an increase from the early hourly products, peak around the mid-day products, and slightly decrease for the afternoon products. These patterns align with the empirical analysis in Section 1.1, particularly the observations illustrated in Figure 2(a) for the buy-side order book (LOB).
@@
Figure 5: Empirical and fitted intensity for Buy/Sell Market order. We estimate the empirical intensity by counting the average number of events during $t,t+\delta t]$, where $N(t)$ the realized counting process for each side of the LOB.
@@
Table 1: Medians ($\pm$ standard deviations) of the estimated parameters of the exponential Hawkes kernel $t\mapsto\alpha e^{-\beta t}$ for Buy and Sell market orders, conditional on delivery hour. Reported are $\alpha$ and $\beta$ (in $\mathrm{s}^{-1}$), the branching ratio $\alpha/\beta$, and the half-life $\ln(2)/\beta$ (in seconds). For intuition, $\alpha=0.12\,\mathrm{s}^{-1}$ corresponds to an instantaneous increase of about $432\,\mathrm{h}^{-1}$ in the conditional intensity after one buy arrival, while $\beta=0.40\,\mathrm{s}^{-1}$ implies a half-life of $\ln 2/0.40\approx 1.73$ seconds.
@@
Figures 6(a) and 6(b) present QQ-plots of inter-event times for all events on the bid side of the order book, focusing on two selected maturities (19h and 21h). For short inter-event times ($\Delta\tau\leq 300s$), the model provides a good fit. However, the fit deteriorates for longer inter-event times, particularly for the shorter maturity (19h). This discrepancy may be attributed to lower trading activity at the start of the trading session. In some sessions, this inactivity significantly extends the tail of the inter-event time distribution, making it slightly left-skewed. This behavior is not fully captured by the calibrated model, potentially limiting its ability to accurately simulate the distribution of long inter-event times.
(a) $Q$-plot for Buy MOs (Simulated vs Realized) for maturity 21h
(b) $Q$-plot for Buy MOs (Simulated vs Realized) for maturity 19h
@@
So far we have considered only the case where the cross-covariances are equal to zero and whilst the counting process $\left(N_{t}^{+},N_{t}^{-}\right)$ describing the arrival of Buy/Sell market orders respectively may be self-exciting, they are independent and thus not mutually exciting. To introduce the excitation between the two types of orders we consider the parameterization of the conditional intensity function given by the full impact matrix kernel (4). Following the goodness-of-fit analysis we have conducted (See Appendix B.1) in the univariate case we can see that overall the fitted model with the spline approximation of the baseline intensity provides a better fit to the data. Hence as a baseline intensity, we will consider the one given by the spline approximation. We present the calibrated parameters for the bivariate case in the Appendix below 5. Our goal is to test the performance of the two models to cost savings compared to the benchmark models; this will be considered in the following section.
In order to calculate the resilience speed of the transient part of the price we will use the approximation derived in [Chen2019]. The authors derived the resilience speed parameter for a TWAP type of schedule, assuming that the market orderflow follows a self-exciting Hawkes process with exponential kernel $t\mapsto\alpha\mathop{}\!\textnormal{e}^{-\beta t}$. Under this setting they showed that a good approximation of the resilience speed is $\rho\approx\frac{1-\alpha/\beta}{T/2}$, where $[0,T]$ the time period where the trader want to liquidate his position. The approximation shows how the resilience parameter links to the self-exciting dynamics of market order flow. As the authors mention this resilience speed compromises only an approximation since different liquidating strategies would in principle lead to different resilience speed parameters $\rho$. To this end we follow this approach and we leave out the calibration of a propagator type of model for future research.
@@ 3.2 Execution Costs
We will now present a summary of the cost estimates for the different liquidation strategies we are going to propose in the following section. In order to perform such a transaction cost analysis we should first find some appropriate cost estimates. Since we do not have any endogenous data, we proceed to estimate various market impact factors from public data disseminated by the exchange and the LOB [EPEXSpotSE]. Our calibration methodology is similar in nature to the analysis presented in [cartea2015algorithmic, glas2020intraday] where an approximation of the half-spread and instantaneous trading costs is studied using LOB market data.
@@ 3.2.1 Instantaneous impact estimation from stochastic order book
Thus, the LOB state at any given time can be fully described by the random vector $\{(M_{i},\Delta P_{i})\}_{i=0,\dots,K_{0}}$. Compared with simple snapshot averages, a non-parametric estimation of the joint distribution $\{(M_{i},\Delta P_{i})\}$ can capture more nuanced stochastic behavior, including asymmetries and heavier tails, leading to better estimates of transaction costs.
@@
$C(\mathcal{U})=\frac{1}{\mathcal{U}}\,\mathbb{E}\bigg[$ $\sum_{j=1}^{K}\mathbf{1}_{\{m_{j-1}<\mathcal{U}\leq m_{j}\}}\Big(\sum_{i=1}^{j}M_{i-1}\,\Delta P_{i-1}$ (25)
$+(\mathcal{U}-m_{j})\,\Delta P_{j}\Big)\bigg],$
where $\mathcal{U}$ is the size of a large market order. In general, estimating such a high-dimensional joint distribution is prone to the curse of dimensionality, requiring a balance between bias and variance. To alleviate complexity, a modified cost function can be used to reduce dimensionality while still offering a meaningful estimate of the instantaneous impact.
@@
$\hat{C}(\mathcal{U})$ $:=\!\frac{1}{\mathcal{U}}\sum_{j=1}^{K}\mathbb{E}\Bigl[M_{j-1}\Delta P_{j-1}\!+\!(\mathcal{U}-\!m_{j})\Delta P_{j}\Big|V_{i,j}\Bigr]$
$\qquad\times\mathbb{P}\big(m_{j}<\mathcal{U}\leq m_{j+1}\big),$ (26)
where $V_{i,j}=\{\omega\in\Omega\colon m_{j}<\mathcal{U}\leq m_{j+1}\}$ and $\nu$ is the market order size, $\Delta P_{j}$ are the price offsets, and $M_{j}$ denotes the liquidity at offset $\Delta P_{j}$. The index $m_{j}$ is the aggregate volume up to level $j$.
@@
To approximate the joint density of the random variables $x=m_{i}$ and $y=m_{i+1}$, we use a bivariate kernel density estimator. Suppose we have $\bigl\{(x_{t},y_{t})\bigr\}_{t=1}^{n}$, where $n$ denotes the number of available data points, drawn from the distribution of $(m_{i},m_{i+1})$. The estimator is:
@@
To estimate the instantaneous market impact, we shall simulate the order book in its stationary state, followed by the submission of orders of size $\mathcal{U}$. For each order, we compute the per-unit cost (price change per MWh) $\hat{C}(\mathcal{U})$ from (3.2.1). Next, we fit a linear relationship:
$\hat{C}(\mathcal{U})=\eta\,\mathcal{U}+\varepsilon,\quad\varepsilon\sim\mathcal{N}\bigl(0,\sigma^{2}\bigr),$
where $\hat{C}(\mathcal{U})$ is the price change per MWh after executing an order of size $\mathcal{U}$ MWh. We interpret $\eta$ (price/MWh2) as the instantaneous market impact factor.
@@
Figure 7 shows the average piecewise-interpolated function describing the linear execution cost coefficient. The median calibrated $\hat{\eta}(t)$ for each hourly contract clearly varies over time; specifically, $\hat{\eta}(t)$ decreases after the trading session begins (measured relative to the last eight trading hours) and slightly increases in the final hour (local trading regime). This aligns with the notion that liquidity diminishes near gate closure, causing higher transaction costs [Bal22, glas2020intraday, KZ20].
Figure 7: Estimated temporary impact coefficients with respect to gate closure for the sell side of the LOB. These impact coefficients are obtained by simulating a large market order that “walks” through the limit order book, assuming a linear relationship $\hat{C}(\mathcal{U})=\eta\,\mathcal{U}+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$ and $\hat{C}(\mathcal{U})$ denotes the price change per MWh for an order of size $\mathcal{U}$.
@@ 3.2.2 Bid-Ask Spread Transaction Costs
Following [Bal22, glas2020intraday, KZ20], the bid-ask spread (BAS) in intraday energy markets exhibits a clear seasonal pattern. In particular, the average BAS decreases significantly at the outset of the trading session and remains nearly constant for the final eight hours of trading. One hour before maturity, when trading is restricted to the local market, the spread widens sharply. This behavior mirrors the pattern observed for the half-spread, which closely tracks the shape of the temporary impact coefficient $\hat{\eta}(t)$. Both observations suggest that liquidity is notably more limited at market open and after market decoupling compared to just before decoupling [glas2020intraday, KZ20].
Consequently, the cost of immediacy, i.e. submitting a market order instead of a limit order, tends to be substantially higher at the start of trading and during the final trading hour. In contrast, during the intermediate trading period, these costs remain relatively lower (see Figure 7).
@@ 4.1 Scaling and discretization of the optimal strategy
We start by describing the use of the optimal strategy described in Equation (24), which reads
$X_{t}^{*}=-\frac{1+\rho h}{\,1-\epsilon\,}\,D_{t}^{*}$ (29)
$+\frac{m_{1}(2+\rho h)}{2(1-\epsilon)\rho}\Biggl(\!1+\!\frac{\rho h}{2+\rho h}\Bigl[\zeta(\eta h)\!+\mu\rho h\,\omega(\eta h)\Bigr]\Biggr)\kappa_{t}.$
We recall that $\kappa_{t}=\lambda_{t}^{+}-\lambda_{t}^{-}$, is the vector of intensity imbalances and are modeled with one of the aforementioned models using the seasonal versions of the baseline intensity, which has been estimated non-parametrically as described in Section 3. We are then going to consider the univariate and bivariate models separately for the backtesting and transaction cost analysis study using the calibrated parameters obtained (See Table 1 for the univariate case and Table 5 for the bivariate case). In order to use the optimal liquidating strategy we consider a subset $\Theta$ of the time interval $\Theta\subset[T_{b,m},T_{e,m}]\subset[0,T]$ possible made of stopping times and trade for each $t\in\Theta$ the quantity
@@
$+\frac{m_{1}}{2\rho}\left(1+\frac{\rho h}{2+\rho h}\times\left[\zeta h\eta)+\mu\rho\omega(h\eta)\right]\cdot s\kappa_{t}\right),$
so that equation (29) holds for $t+\Delta t$ if $s=1$. In order to calculate the resilience part of the execution price we consider the simplified expression
$D_{t}^{*}=\sum_{\tau\leq t}\Delta N_{\tau}\left[G(t-\tau)-G(\infty)\right].$ (31)
The propagator kernel is chosen as in the OW framework $G(u)=(1-\mu)\exp(-\rho u)+\mu$ and $\Delta N_{\tau}$ the jumps of the order-flow. As we have already described we have a positive (negative) unit jump when we have an arrival of a buy (sell) market order. We also introduce the scaling factor $s\in[0,1]$ that multiplies both $\kappa_{t}$ and $D_{t}$ and is a hyper-parameter that tunes the leveraging effect of the strategy, describing the deviation of the whole strategy from the standard liquidation scheme in [Obizhaeva2005]. We can observe that in the case s=0s\!=\!0 we derive the static strategy, which is nothing else by the block trades that have been derived in the OW strategy, and in the case s=1s\!=\!1 we derive the strategy given by optimal execution strategy we derived under the Marked Hawkes process which drives the order flow and presented in Theorem 2.2.